OpenStudy (anonymous):

HELP nid it today please answer!! 5x - 3y = 15 ( ,5),( ,0),(2, ) x + 4y = -2 (0, ),( ,0),(1 , )?

OpenStudy (anonymous):

i got your back

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

\[5x - 3y = 15\\ ( ,5),( ,0),(2, ) \]

OpenStudy (anonymous):

\[5x-5\times 5=15\] \[5x-25=15\] \[5x=40\] \[x=8\]\[(8,5)\]